Can you please help me solve this one...
Find the x and y intercepts of this equation: -7x + -3y = -21

Answer by mananth(16946)

-7x + -3y = -21
y Intercept
Plug x=0
0-3y=-21
/3
y=7 the y intercept.
X intercept
Plug y=0
-7x+0=-21
/-7
x=3 the X intercept